OpenStudy (anonymous):

What is the value for x in the proportion x\x+10=7\12

Parth (parthkohli):

\(\color{Black}{\Rightarrow \Large {a \over b} = {x \over y} \Longrightarrow ay = bx }\)

OpenStudy (lgbasallote):

\[\frac{x}{x+10} = \frac{7}{12}\] cross multiply

Parth (parthkohli):

\(\color{Black}{\Rightarrow 7(x + 10) = 12x }\)

Parth (parthkohli):

Distribute first.

OpenStudy (lgbasallote):

\[\LARGE \Rightarrow a(b \pm c) = ab \pm ac\]

OpenStudy (anonymous):

when i cross multily I got 12x\70x

Parth (parthkohli):

How so? o.O

OpenStudy (lgbasallote):